### v2

Abstract
In this study, we address the challenge of detecting “discursive communities” on X/Twitter by focusing on the role of verified users as the main content creators in online political debates. The analysis centers on three major Italian political events in 2022 – the Presidential election, a governmental crisis, and the general elections – occurring before the introduction of paid account verification. We propose and compare two novel methodologies, MonoDC and BiDC, which exploit, respectively, the retweet network among users and a similarity network based on shared audiences, while integrating a maximum entropy null model to filter out the inherent noise in online social networks. Our results demonstrate that leveraging verified users—considered as indicators of prestige and authority—leads to significantly clear community partitions that closely reflect the actual political affiliations, outperforming standard community detection algorithms applied to the entire retweet network. Moreover, the comparison of different methodologies and user sets suggests that the status conferred by the blue verification tick plays a dominant role in shaping online discourse, with important implications for platform governance, especially in light of the recent shift to paid verification.
1 Introduction
Online communities naturally emerge from users’ tendency to follow and interact with others who share similar interests, experiences, connections, or political views [1, 2, 3]. Such communities can provide valuable support and shared knowledge [4, 5, 6]; however, they can also form “echo chambers,” where members are exposed only to a narrow range of perspectives that reinforce their existing beliefs [7, 8, 9, 10, 11, 12, 13]. To fully understand the influence of user communities on information diffusion and opinion dynamics, it is crucial to develop effective methods for detecting these communities and capturing their key structural features.
In the literature, several refined approaches have been proposed to enhance standard community detection methods by tailoring them to the context of online social networks. For example, Citraro and Rossetti [14] incorporate account attributes to strengthen the conventional definition of communities. Other approaches leverage the varied roles played by different accounts in online debates. Although Twitter/X was initially designed as a many-to-many platform, numerous studies have observed that it actually functions as a few-to-many network, where a small number of influential content creators generate the majority of posts [15, 16, 17]. In this context, it becomes natural to focus on the leading content creators for detecting communities on Twitter/X.
Identifying which accounts should be considered as leading content creators is, therefore, essential. In [4], influential users—defined as those with more followers than followees and with more mentions received than given—are identified as the primary figures in the debate. Similar approaches, based on user activity, have also been explored [15, 16, 17]. Other studies place more emphasis on account prestige than on activity, considering verified users as the leading content creators [18, 19, 20].
Verified users on X/Twitter have an interesting history: they were first introduced in June 2009 to curb the impersonation of notable persons or organizations [21, 22]. In this early stage, the platform assumed responsibility for identifying key public figures and verifying their identities with a blue checkmark, while standard users were not part of the verification process [23]. Efforts to broaden the pool of verified accounts through user applications were unsuccessful [21]. Eventually, in November 2022, verification became a feature of X’s Premium accounts, available upon payment [24]—a change that has sparked concerns regarding the influence of verified users on online debates [25, 26].
In the present study, we investigate the method proposed in [18] using three distinct datasets collected during 2022, in a period before blue ticks could be purchased via X’s Premium subscription. Our framework comprises: (i) identifying verified users as the leading content creators; (ii) constructing an informative network of verified users by filtering out random noise through a comparison between empirical data and an appropriate null model; (iii) detecting communities within this sub-network; and (iv) propagating the labels to all other users. Following the terminology introduced by [27], the resulting partition is referred to as “discursive” communities.
A key component of our approach, in addition to leveraging verified users, is the elimination of random noise from the system by comparing empirical data against a maximum entropy statistical benchmark. As numerous studies have shown, removing random noise is crucial in online social networks to extract the relevant signal regarding users’ preferences and behaviors [18, 28, 29]. In this regard, maximum entropy null models provide an ideal statistical benchmark, being maximally random apart from the constraints imposed [30].
Figure 1: MonoDC’s pipeline. Starting from the data from Twitter/X, we build a bipartite directed network between accounts and posts; a directed arrow from a user to a post represents authorship, while in the opposite direction represents a retweet. Then, the content of the bipartite network is projected into a monopartite directed network of users. Such a network is further statistically validated using the expectations of a maximum entropy null model for bipartite directed networks. Then, the sub-network relative to content creators is analysed, and communities are detected using standard algorithms. Finally, the so-obtained labels are propagated to standard users through a label propagation algorithm.
Compared to previous work, we distinguish between two main versions of the algorithm that differ in how step (ii) is implemented. In $\mathrm{MonoDC}$ (Monopartite-network-induced Discursive Communities), we consider the direct interactions —i.e., retweets— among content creators; a cartoon representation of the various steps of the algorithm is displayed in Fig. 1. In $\mathrm{BiDC}$ (Bipartite-network-induced Discursive Communities), we instead construct a similarity network based on the shared audience between pairs of leading content creators; a cartoon representation of the various steps of the algorithm is displayed in Fig. 2. Additionally, we explore various classes of leading content creators beyond verified users.
Our proposed methods were tested on three large datasets linked to the primary political events in Italy during 2022, before user verification became a feature of Premium accounts. To validate our results, we manually annotated approximately 300 verified accounts to determine their political leanings and compared different clustering algorithms to assess their ability to detect this signal.
The key finding is that both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$, when using verified users as the main content creators, significantly outperform standard algorithms. They produce clearer partitions that consist of a few meaningful communities, which accurately reflect the political landscape during the studied period. Specifically, $\mathrm{MonoDC}$ appears to capture political parties more effectively—likely because politicians predominantly retweet their party leaders—while $\mathrm{BiDC}$ better identifies political alliances due to the shared construction of narratives. Importantly, our results deteriorate markedly when the set of main content creators is altered.
Figure 2: BiDC’s pipeline. Starting from the data from Twitter/X, we build a bipartite undirected network between the accounts of the leading content creators and the one of standard users; a link is present if the standard users retweeted the given content creator at least once. Then, the information of the bipartite network is projected into a monopartite network of leading content creators. Such a network is further statistically validated using the expectations of a maximum entropy null model for bipartite undirected networks. Finally, the so-obtained labels are propagated to standard users through a label propagation algorithm.
In summary, our study demonstrates that verified accounts play a central role in shaping discursive communities on X/Twitter. This evidence carries important implications, especially considering that our datasets are limited to the period before the implementation of X’s Premium accounts, when blue checkmarks were granted based on the platform’s autonomous verification of key accounts. While we cannot conclusively determine whether the effectiveness of verified users as leading content creators stems solely from the prestige of the verification or from other factors, our findings indicate that definitions based exclusively on user activity yield inferior performance. If the prominence of verified users in steering the online discourse is indeed due to the prestige associated with blue ticks, then introducing verification as a paid feature could pose serious challenges to the platform’s integrity, potentially enabling malevolent actors to purchase authority to distort the political agenda.
The paper is organised as follows. In Section 2, we introduce the methodology adopted in this study. Specifically, we present maximum entropy null models and their application to the detection of discursive communities (Subsections 2.1 and 2.2). We also define the V-measure [31], the metric used to compare different partitions (Subsection 2.3). Section 3 presents our results. We begin by illustrating the limitations of standard methods in reproducing non-trivial benchmarks observed in real-world data (Subsection 3.3). We then show how $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ address these issues when verified users are identified as the primary content creators (Subsection 3.4). Additionally, we show that the performance of both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ deteriorates when different groups of leading content creators are considered (Subsection 3.5). Finally, in Section 4, we summarise our findings and discuss their potential implications.
2 Methods
2.1 Maximum-entropy null models and discursive communities
As underlined by many studies across a multitude of applications  [18, 19, 20, 27, 32, 33, 34, 35, 36, 28, 29, 13], using a proper statistical validation is critical for the analysis of online social networks (OSNs) that are particularly prone to noise. Entropy-based null models [30] represent a tailored and, at the same time, unbiased tool for detecting the statistically significant signal in complex networks. The method essentially consists in: i) defining an ensemble of graphs, i.e. the set of all possible graphs having the same number of nodes as the observed network; ii) performing a constrained maximization of the Shannon entropy of the ensemble, fixing some constitutive properties of the original system [37]; iii) maximizing the (log-)likelihood of the observed network [38], to get the numerical values of the parameters of an effective fitness model.Remarkably, this framework mimics Jaynes’ derivation of statistical ensembles in Physics starting from Information Theory [39].
Let us start from a real network $G^{*}$ (in the following, all quantities related to the observed network will be denoted by an asterisk *). Firstly, let us define a set $\mathcal{G}$ – the ensemble– including all possible graphs having the same number of nodes $N$ as in $G^{*}$ [^1] Since the number of nodes is fixed in the randomization, we drop the asterisk * from $N$.. Let us identify a vector of quantities $\vec{C}$ that are fundamental for the description of the real system. A proper benchmark for $G^{*}$ would be completely random, but for the information contained in $\vec{C}(G^{*})$: in this way it is possible to highlight all properties of $G^{*}$ that are not trivially due to $\vec{C}$. Translating this rationale in maths, we need to perform a constrained maximization of the Shannon entropy (to have a maximally random benchmark) associated to the ensemble, where the constraints are exactly $\vec{C}(G^{*})$. In this sense, the (ensemble) average $\langle\vec{C}\rangle$ will be set exactly to $\vec{C}(G^{*})$.
If $S$ is the Shannon entropy, defined as
$S=-\sum_{G\in\mathcal{G}}P(G)\ln P(G),$
then we can perform the constrained maximisation using the method of Lagrange multipliers, i.e. by maximising $S^{\prime}$ defined as
$S^{\prime}=S+\vec{\theta}\cdot\Big{(}\vec{C}(G^{*})-\langle\vec{C}\rangle\Big{)}+\alpha\Big{(}1-\sum_{G\in\mathcal{G}}P(G)\Big{)},$
where $\vec{\theta}$ is the vector of Lagrange multipliers associated with $\vec{C}$ and $\alpha$ is the Lagrangian multiplier associated to the normalization of the probability $P(G)$. The probability solving the maximization problem above has the standard functional form of exponential random graphs:
$P(G)=\frac{e^{-\vec{\theta}\cdot\vec{C}(G)}}{Z(\vec{\theta})},$ (1)
where $Z=e^{\alpha+1}=\sum_{G\in\mathcal{G}}e^{-\vec{\theta}\cdot\vec{C}(G)}$ is the partition function. In order to get the numerical value of the Lagrangian multipliers $\vec{\theta}$ we have to explicitly impose
$\langle\vec{C}\rangle=\vec{C}(G^{*}).$ (2)
It can be shown that the condition in Eq. (2) can also be obtained through the maximization of the (log-)likelihood [38].
The framework described above is particularly flexible and can be adapted to different kinds of networks. In the present paper, we are interested in its bipartite versions for bipartite undirected and bipartite directed networks, i.e. respectively, the Bipartite Configuration Model (or BiCM [40]) and the Bipartite Directed Configuration Model (or BiDCM [41, 18]). In both cases, the constraints imposed in the entropy maximisation are, respectively, the undirected and directed degree sequences. More details about BiCM and BiDCM can be found in Appendices A.1 and A.2.
2.2 $\mathrm{MonoDC}$ and $\mathrm{BiDC}$
In the following, the two methods that we are proposing are presented. They both leverage the presence of leading content creators, i.e. users mostly posting original messages and less inclined to repost the messages created by others. There are multiple ways to define leading content creators: at the present level, we prefer to remain as general as possible. Therefore, in the following those users are called $\top$-⊤ -users, while the $\top$ users are called $\bot$-⊥ -users.
2.2.1 $\mathrm{MonoDC}_{\top}$
At least in the Italian political debate, it is common to observe that less relevant political figures are particularly active in liking and reposting the content created by more central leaders [^2] https://www.ilpost.it/2024/10/31/post-giorgia-meloni-social-commenti-parlamentari-fdi/. There could be multiple reasons for such a behaviour: intercept the audience of the leaders, demonstrate loyalty to either the leaders, the party or the objective of the campaign, or simply to contribute to the campaign of the leaders by increasing the visibility of their content. Therefore, it is reasonable to expect the activity of politicians to be much more focused on the activity of their party than other users.
To capture such a phenomenon, the MonoDC studies the interactions among content creators through their activity in creating new contents or reposting the one created by others. The steps of MonoDC are represented as a cartoon in Fig. 1. We describe data from Twitter/X as a bipartite directed network between users and posts in which edges from users to posts indicate authorships, while the vice versa indicates retweets. The information contained in the directed bipartite network is then projected to describe the retweet network, i.e. a directed multilink network in which nodes are users and the number of edges from user $A$ to user $B$ is the number of retweets that $B$ made to the post written by $A$. In this representation, we expect the (bipartite directed) degrees to play a role, since they represent the activity of users in either creating new posts or retweeting existing ones (represented, respectively, as the out- and in-degree in the bipartite directed network of the user nodes) and the virality of the various posts (i.e. the out-degree of the post nodes). Therefore, we can use a maximum entropy null model for directed bipartite networks as a statistical benchmark to validate the empirical retweet network, i.e. the Bipartite Directed Configuration Model (BiDCM) mentioned in Section 2.1. Technical details about BiDCM can be found in Appendix A.2.
The result of such validation is a directed binary network, in which an edge represents that the retweet activity between the two nodes is non-trivial, i.e. it cannot be explained by just the attitude of users towards posting and reposting and the virality of the messages shared. We further restrict the validated network to only users in the set of content creators, i.e. the $\top$ user, in our jargon.
The main idea is that the null model is not informed about the orientation of the users and therefore expects the debate to be “flat”, i.e. without any particular focus of the users towards the content created by some of the accounts. In this sense, communities should emerge naturally by comparing empirical data with the null model. Such a structure can be further captured by running a community detection algorithm on the validated retweet network.
Finally, the labels obtained by the community detection are extended to $\bot$ users (i.e. the users that are not in the $\top$ set) by implementing Raghavan’s Label propagation algorithm [42] on the retweet network, using as label seeds the ones obtained by the community detection run on the validated retweet network restricted to $\top$-users.
2.2.2 $\mathrm{BiDC}_{\top}$
The intuition behind $\mathrm{BiDC}_{\top}$ is to capture discursive communities through the perception of “standard” users of the similarity in the content created by the leaders of the debates [18, 43, 27, 19]. Such a phenomenon is more stressed during political debates, being more or less intense depending on the electoral system of the country [43].
Following the procedure outlined in Fig. 2, we built the bipartite network between leading content creators and “standard” users: a link is present if the standard user has retweeted at least once the corresponding content creator. We then projected the bipartite network to a monopartite co-occurrences network among content creators. As in the case of $\mathrm{MonoDC}_{\top}$, we validate the network by comparing the empirical data with a maximum entropy null model for bipartite networks, i.e the Bipartite Configuration Model (BiCM) mentioned in Section 2.1. In this case, the information about the degree of the users is capturing the number of different retweeters and the number of different content creators retweeted for $\top$-⊤ - and $\bot$-⊥ -users, respectively; more details about the implemented maximum entropy null model can be found in Appendix A.1.
BiCM, while accounting for the degree sequences of both layers, is not informed about the presence of communities and therefore expects the co-occurrence network not to have a modular structure. In this sense, any non-trivial focus of the users towards groups of content creators should be captured by comparing empirical data against the null model. Such a structure can be further refined by running a community detection algorithm on the validated projection network.
Similarly to $\mathrm{MonoDC}_{\top}$, labels are extended to $\bot$ users by using Raghavan’s Label propagation algorithm [42] on the retweet network, using as label seeds the ones obtained by the community detection run on the validated retweet network.
2.3 Measuring the similarity of two partitions
To assess the similarity between two ways of partitioning the same set of users, we use of the V-measure ($\mathrm{VM}_{\beta}$), a parametric measure of cluster similarity proposed by [31].
Let $S(P)$ be the Shannon entropy of $P$, and $S(P\mid Q)$ be the conditional entropy of $P$ given $Q$. If $P$ and $Q$ are two partitions, the $\mathrm{VM}_{\beta}(P,Q)$ is defined as:
$\mathrm{VM}_{\beta}(P,Q)=\frac{(1+\beta)I(P,Q)}{S(P)+\beta S(Q)}$
where $I(P,Q)$ is the mutual information between $P$ and $Q$.
For $\beta=1$ we have
$\mathrm{VM}_{1}(P,Q)=\frac{2I(P,Q)}{S(P)+S(Q)}=1-\frac{VI(P,Q)}{S(P)+S(Q)}$
i.e., the $\mathrm{VM}_{1}$ accounts for the mutual information between the two partitions in a way that does not depend on their entropy, being equivalent to the Variation of Information [44], normalized by the total entropy of the partitions. With respect to other information-theoretic metrics, the $\mathrm{VM}_{1}$ and the VI have the advantage of being two distances – in particular, of satisfying the triangle inequality. Further, the $\mathrm{VM}_{1}$ always takes values in $[0,1]$, making it possible to directly compare the scores obtained for different combinations of algorithms and/or datasets. By letting $\beta$ vary in $[0,+\infty)$, we consider both scenarios in which one partition consists of sub-communities (e.g., political movements/trends within parties) or super-communities (e.g., political coalitions) of the other one. More details about the V-measure can be found in Appendix B.
3 Results
3.1 Data and networks
To compare our approach for discursive community with standard community detection algorithms, we collected Twitter data regarding the three main political events debated on Italian Twitter in 2022. For each such topic of discussion, we identified a set of relevant keywords that were used to collect tweets using Twitter’s Academic API. Let us remark that data were downloaded before user’s verification became a feature of the X’s Premium subscription [^3] https://en.wikipedia.org/wiki/Twitter_verification.. A summary of the datasets considered in this paper is reported in Table 1.
Table 1: A summary of the datasets considered in the present paper
Dataset Unique users Unique tweets Annotated users Verified users Influential users H-$\geq$3 users
President1 119.018 711.497 292 1148 5804 1296
Crisis2 113.876 1.414.181 229 1004 7038 2554
Elections3 222.356 2.784.951 239 1555 12184 4382
  • 1
    Debate around the election of the President of the Republic. The political debate ended with the re-election of Sergio Mattarella on January 29, 2022.
  • 2
    Debate around the political crisis that led to the fall of the Draghi government on July 21, 2022.
  • 3
    Debate around the Italian general elections held on September 25, 2022, that led to Giorgia Meloni becoming the new Prime Minister.
3.2 The benchmark
In general, there is no reason to expect data to cluster due to some metadata [45]. Multiple effects can cause –especially in online social networks– different users to interact with the same content creators. Nevertheless, due to the strictly political nature of the debates captured by our data, we expect to observe a signal of the political orientation of the users in the cluster that can be used as a benchmark for the classification methods under study.
To construct our benchmark labeling, we first focused on just the verified accounts of Italian politicians and political parties. We choose verified accounts since their identities are certified by the platform. We manually labeled these accounts according to their political party and then grouped these labels according to the coalition each party belonged to during the general elections. This procedure yielded two annotation levels, which we refer to as “parties” and “coalitions”.
To check if this benchmark is well-posed, we studied the retweet network restricted to the annotated users. Retweets are usually considered as a sort of endorsement of the content shared [1], so we expected users with a similar political orientation to cluster in this network. The modularity of the partitions induced by the annotations, whose values are all above 0.6 (cfr. Table 2), confirms our intuition and supports the idea of measuring the quality of a community detection algorithm also based on its ability to retrieve our manual annotations.
Table 2: Retweet networks of annotated users and modularity (Q) of the partition induced by the manual annotation at the party and coalition level.
Dataset Nodes Links $Q(\text{parties})$ $Q(\text{coalitions})$
President 169 219 0.763 0.618
Crisis 145 246 0.721 0.641
Elections 154 355 0.777 0.649
To compare different algorithms not just based on the small minority of users that we manually annotated, we extended the annotations to the rest of the network using a well-known Label propagation algorithm [42] –taking the most frequent label out of 100 independent realizations. These propagated labels cannot be considered equally accurate, but they arguably represent the best possible guess of a user’s political leaning, which can be made only based on how that user is connected to verified politicians in the retweet network.
Let us notice that such an approach favors our models, in principle, since we are using the same label propagation algorithm to label standard users. In a sense, if we correctly tag $\top$ users, we will implicitly also meet the benchmark for $\bot$ users. It is indeed true, but it is an honest way to propagate the labels of verified accounts, as we cannot trust, in general, the one of other accounts that could be present in the user description.
3.3 Limits of standard methods
Let us start by considering the partitions obtained running four off-the-shelf community-detection algorithms on the entire retweet network: (i) Louvain, a heuristic for modularity maximization; (ii) Label propagation, a message passing algorithm where each node tries to align with the majority of its neighbors; (iii) Infomap, which is based on finding the optimal encoding for random walks on the graph; (iv) Inference, whose rationale is inferring the parameters of a stochastic block-model through a maximum-likelihood approach.
For Louvain, Label propagation and Infomap, we used the implementation available in igraph (igraph.org, Python interface). For Inference, we used the minimize_blockmodel_dl() function provided in graph-tool (graph-tool.skewed.de), that fits a stochastic block-model to the given graph by minimizing the model’s description length. To guide the inference process, we set it explicitly to search for a planted-partition, that is, exclusively for assortative communities. In addition, we refined the algorithm as recommended in graph-tool’s documentation, that is, by running 100 times the multiflip_mcmc_sweep() method with zero temperature and 100 iterations.
We ran each algorithm 100 times [^4] Each single run of the Louvain algorithm is already obtained taking the partition with the greatest modularity out of 100 independent executions with shuffled vertices. Similarly, each run of the Label propagation algorithm is already obtained by taking, for each node, the most frequent label out of 100 independent executions. and measured: the $\mathrm{VM}_{1}$ between each partition, restricted to the annotated users, and the annotations, both at the party and coalition level; the pairwise $\mathrm{VM}_{1}$ between partitions obtained with different algorithms, restricted to the verified users. Fig. 3 reports these $\mathrm{VM}_{1}$ values, considering only the parties annotations. We focused on party annotations, but similar results are observed with other algorithms and with coalitions.
Fig. 3 highlights the limited accuracy in recognizing the affiliation of renowned politicians –$\mathrm{VM}_{1}\in[0.4,0.6]$– of the communities assigned to verified users by different off-the-shelf algorithms.
Furthermore, as shown in Fig. 10 in Appendix C, standard algorithms show great variability of the communities assigned to verified users by different off-the-shelf algorithms –pairwise $\mathrm{VM}_{1}\in[0.3,0.7]$. Overall, these findings motivate the research for a different approach to the detection of discursive communities on Twitter/X.
Figure 3: Limits of standard methods. The $\mathrm{VM}_{1}$ between partitions obtained with off-the-shelf algorithms, restricted to the annotated users, and the parties’ annotations. Standard methods show a limited accuracy in recognising the affiliation of renowned politicians.
3.4 $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$
We now show the results of our methods, i.e. $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, when we choose verified users as the class of content creators: in this case, we will write $\top=V$. In the following, the communities on the network of $V$ users are obtained by running the Louvain algorithm on the projected network and are then propagated to the rest of the network. We tested our community detection algorithm, as the Label propagation and Infomap: their results are very similar to those obtained with Louvain ($\mathrm{VM}_{1}>0.8$) and yield similar results also for both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$. Communities obtained with the inference-based method are instead way less accurate in reconstructing the annotations.
3.4.1 Comparison with standard methods
In Fig. 4, we use the $\mathrm{VM}_{1}$ to show how accurately each method reconstructs the manual annotations at the party and coalition level, respectively in the first and second row. Each violin represents the distribution over 100 independent runs, and we compare $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with the partitions obtained running the Louvain algorithm on the entire retweet network –which performs no worse than other algorithms (cfr. Fig. 3). $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ systematically outperform Louvain, with $\mathrm{MonoDC}_{V}$ generally preferable in identifying parties while $\mathrm{BiDC}_{V}$ in identifying coalitions. The fact that in all three cases the community detection algorithm used is Louvain signifies that what ultimately matters is focusing on the right network.
For completeness, the comparison between the performances of our methods and Louvain is presented in Fig. 11 in Appendix C. Even in this case, both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ outperform Louvain. Nonetheless, let us note that, when comparing the results of the propagated labels of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}$ with the annotations of verified users propagated using the same label propagation algorithm, our methods have an advantage. Nevertheless, the performances of standard methods are worse, even when limited to annotated users.
Figure 4: Comparison of the performances of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with standard methods. The $\mathrm{VM}_{1}$ between partitions of the annotated users, obtained with different methods, and the annotations at party (first row) and coalition (second row) level. $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ systematically outperform Louvain, with $\mathrm{MonoDC}_{V}$ generally preferable in identifying parties while $\mathrm{BiDC}_{V}$ in identifying coalitions.
It could be argued that standard algorithms, when executed on the entire retweet network, return sub- or super-groups of the political parties that $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ cannot detect. Fig. 5 shows that this is not the case: when compared with manually annotated parties (top) and coalitions (bottom), $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ exhibit a consistently greater average $\mathrm{VM}_{\beta}$ than standard methods even for $\beta\neq 1$, i.e. at essentially all scales at which the political spectrum can be observed.
Figure 5: Comparison of the performances of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with standard methods at different resolution levels. The $\mathrm{VM}_{\beta}$ between partitions of the annotated users, obtained with different methods, and the annotations, for different values of $\beta$ and both at the parties (top) and coalitions (bottom) level. Each point is the average over 100 runs of the method indicated by the colour and marker. Analogous results are obtained when considering the propagated annotations. When compared with manually annotated parties (top) and coalitions (bottom), $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ exhibit a consistently greater average $\mathrm{VM}_{\beta}$ than standard methods even for $\beta\neq 1$, i.e. at essentially all scales at which the political spectrum can be observed. In this sense, standard algorithms, when executed on the entire retweet network, do not return sub- or super-groups of the political parties that $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ cannot detect.
In Fig. 6, the Sankey diagrams corresponding to each evaluated method provide a detailed depiction of the political composition of the extracted communities. $\mathrm{MonoDC}_{V}$ exhibits remarkable precision in delineating party structures —successfully distinguishing individual parties such as Azione and Italia Viva, which presented a unified list during the elections— and further reveals latent political currents within party boundaries. Conversely, $\mathrm{BiDC}_{V}$ proves highly effective in reconstructing the official electoral coalitions, achieving a near-exact correspondence with their institutional configurations. The Louvain algorithm, applied to the entire retweet network, yields a substantially noisier community structure that conflates partisan and coalition-level affiliations. Notably, Forza Italia emerges as an isolated community, distinct from its allied parties, while several political figures are clustered within communities dominated by their ideological adversaries.
(a) Louvain on RT network.
(b) $\mathrm{MonoDC}_{V}$.
(c) $\mathrm{BiDC}_{V}$.
Figure 6: Qualitative comparison of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ with Louvain on the Elections dataset. For each of the considered methods, the Sankey diagrams show the composition of the extracted communities in terms of political parties. $\mathrm{MonoDC}_{V}$ very accurately detects parties (Azione and Italia Viva ran together as a single list at the elections), plus some political movements/trends within parties; $\mathrm{BiDC}_{V}$ almost perfectly detects the official political coalitions of those elections; Louvain, executed on the entire retweet network, provides a more noisy classification that is a mix of the two levels, with Forza Italia separated from its coalitions and many politicians contributing to communities led by some of their rivals.
Finally, since $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ capture the political leaning of users at different scales, in Fig. 7 we use the $\mathrm{VM}_{\beta}$ to check whether the clusters found by $\mathrm{MonoDC}_{V}$ are subclusters of those found by $\mathrm{BiDC}_{V}$. This seems to be true, at least in part, and especially for the two datasets, Crisis and Elections, which refer to a political phase in which the coalitions were more clearly defined.
Figure 7: Different resolution of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$. The $\mathrm{VM}_{\beta}$, for variable $\beta$, between partitions of all users obtained running $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$. The clusters found by $\mathrm{MonoDC}_{V}$ are subclusters of those found by $\mathrm{BiDC}_{V}$, at least in part, and especially for the two datasets, Crisis and Elections, which refer to a political phase in which the coalitions were more clearly defined.
3.4.2 Final remarks
Both $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ benefit from a pre-processing stage capable of extracting, from the original retweet data, a more informative network of just such leading content creators. The role that statistical validation plays in the two cases is, however, incomparable.
In $\mathrm{MonoDC}_{V}$, the validated retweet networks of verified accounts of politicians include 94-97% –depending on the dataset– of all edges present before the validation. The validation is performed with respect to the patterns emerged from the analysis of the entire network and the fact that almost all retweets made by politicians are statistically significant means that their activity on X is extremely focused on just the content produced by a limited niche of users. Since these users are their party’s leaders and fellow members, $\mathrm{MonoDC}_{V}$ is the approach that best captures the loyalty to the party and, hence, the difference between parties that apparently occupy a similar position in the political spectrum. The price to pay for this finer classification is its minor stability, as can be seen in Fig. 4. In this sense, $\mathrm{MonoDC}_{V}$ seems particularly promising to examine the mesoscale structure of such a subgraph to infer the main poles of the political discussion, since the phenomenon we are leveraging on is particularly strong.
Differently from the $\mathrm{MonoDC}_{V}$’s case, for $\mathrm{BiDC}_{V}$ the impact of the statistical validation is critical: the number of links in the validated network is 7-12% –depending on the dataset– of those present in the non-validated projection. When the political sides are clear, this validated signal coming from the audience returns a very stable partition (cfr. Fig. 4). In addition, $\mathrm{BiDC}_{\top}$ has the advantage that the similarity between any two content creators $u,v\in\top$ can be established even when $u$ and $v$ did not directly interact, making $\mathrm{BiDC}_{\top}$ more flexible, robust and widely usable than $\mathrm{MonoDC}_{\top}$.
For the sake of completeness, we ran the same set of community-detection algorithms considered in Section 3.3 on the non-validated network, as well as on the validated projections with different significance levels [^5] To properly target the significance level in validating the projected network, we implemented the False Discovery Rate (FDR) procedure [46], see Appendices A.1.1 and A.2.1., cfr. Fig. 12 in Appendix C. The results on the non-validated network are very poor, whereas the significance level does not seem to have a great impact on the inferred similarity network.
3.5 Other classes of content creators
In the previous subsections, we examined the behaviours of the $\mathrm{MonoDC}$ and $\mathrm{BiDC}$ after identifying verified users as relevant content creators, i.e. the main leaders of the shared discourses. Verified users ($V$) are a very natural choice for the $\top$ layer of both $\mathrm{MonoDC}_{\top}$ –because their behavior on the platform is bound to their role outside of it– and $\mathrm{BiDC}_{\top}$ –because they are generally known to and easily categorizable by the audience.
There are other ways, however, to select users based on their actual or presumed influence on online discussions. For instance, [4] defined “influential” users those who have more followers than friends (i.e. the accounts a user follows, sometimes referred to as “followees” in the literature), and who, in the debate under study, received more mentions than those they made. Another way to measure the impact of a user in a community, commonly employed for scientists and academics, is the Hirsch index [47] –shortly, H-index– which can be adapted to X’s mode of operation by saying that a user $u$’s H-index in the debate is the largest $h$ such that at least $h$ tweets authored by $u$ received at least $h$ retweets each.
In the following, we use $I$ to denote the set of influential users and $H$ for the set of top H-index users (defined as users having H-index3absent3$\geq$ 3≥ 3, see Fig. 13 in Appendix C for a justification of the threshold). We aim to assess both what happens when $\top=I$ or $\top=H$ is used in $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, and how stable and accurate is the classification obtained for these two sets of central users when compared with the entire dataset.
(a) $\mathrm{MonoDC}_{\top}$, for different choices of $\top$, compared with annotations at the parties level.
(b) $\mathrm{BiDC}_{\top}$, for different choices of $\top$, compared with annotations at the coalitions level.
Figure 8: Comparison of the partitions obtained with $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, for different choices of $\top$. The $\mathrm{VM}_{1}$ between partitions obtained with different choices of $\top$ and the annotations, for the Elections dataset. For each method, we used the level of annotation that the method more accurately reproduces, that is, parties for $\mathrm{MonoDC}_{\top}$ (a) and coalitions for $\mathrm{BiDC}_{\top}$ (b). The columns refer to the user set upon which the VM is computed. For instance, “user set = influential” means that in the second column we only consider how influential users have been labeled by different methods, temporarily ignoring the rest of the network.
In Fig. 8, the $\mathrm{VM}_{1}$ is used to compare partitions obtained with different choices of $\top$ with the (propagated) annotations ($x$-axis). Here we consider the Elections dataset, but similar results hold for the other datasets.
In Appendix C), we also compare the cases $\top=H$ and $\top=I$ with the case $\top=V$, taken as a reference.
The choice of the $\top$ set has an impact on both $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$. Especially for $\mathrm{BiDC}_{\top}$, $\top=V$ yields the best overlap with the annotations. It could be argued that it is due to $V$ being a small set that includes all manually annotated users, but $H$ outperforms $I$ despite the former includes just 20-50% (depending on the dataset) of the annotated public figures, compared with 80-90% of the latter.
This supports the claim that $V$ users are the key to understanding the sides of the public discourse.
4 Conclusions
In this paper, we focused on X/Twitter and on so-called discursive communities, i.e. the sub-groups that naturally emerge in Online Social Networks (OSNs) as users cluster around shared narratives, language, or value systems during public debates. What we observe in OSNs is a network of interactions, where edges exist for a combination of factors – e.g. the visibility, the status and the behaviour of different users on and off the platform, and the platform’s recommendation system – and where understanding what led to the creation of a community is almost as important as detecting the community itself.
In this paper, we investigate the role of verified users in shaping discursive communities in the pre-Premium accounts era. Introduced in 2009, verified accounts were originally intended to prevent impersonation of public figures or organizations [21, 22]. At first, the platform itself selected which accounts to verify, based on their relevance in public discourse. These accounts were marked by a blue checkmark next to their name. While verification helped authenticate the presence of public figures, it also introduced a hierarchy among users—one that arguably amplified the influence of verified accounts in online debates [48].
Despite these concerns, in November 2022, following Elon Musk’s acquisition of Twitter, verification became a paid feature available through X’s Premium subscription [24].
We conduct our analysis around three key political events in Italy during 2022: the election of the President of the Republic, the political crisis, and the subsequent national elections. Our initial observation is that a strong signal of political alignment is detectable in the data: when clustering verified politicians in the retweet network by political party, the resulting partition shows high modularity. We then explore the limitations of existing community detection algorithms. When comparing their output to manually annotated data, both descriptive and inferential methods often produce noisy and imprecise community structures.
To address this, we evaluate the performance of two algorithms designed to identify discursive communities: $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$, both of which begin by labeling verified users, given their central role in public debate. The $\mathrm{MonoDC}_{V}$ approach is grounded in a typical political dynamic, where less prominent figures tend to retweet content from party leaders. Conversely, the $\mathrm{BiDC}_{V}$ method—originally proposed by [18]—builds an undirected similarity network between content creators by analyzing audience overlap. In both cases, we reduce the effect of random noise—an inherent feature of OSNs [28, 29]—by comparing empirical data to a maximum entropy null model, which accounts for user activity [30].
Both $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}_{V}$ successfully reproduce the organization of political debates into discursive communities shaped by parties or alliances. Notably, using verified users as leading content creators consistently outperforms alternative selections based solely on user activity. This suggests that the status conferred by verification plays a more decisive role in structuring the online conversation than activity alone.
In summary, our work offers two take-home messages. First, in X/Twitter datasets from before November 2022, leveraging verified users allows for a clearer and more accurate identification of discursive communities. Alternative approaches relying solely on user activity result in noisier partitions, suggesting that user status is a major factor in shaping the online debate.
Second, our findings are based on comparisons conducted before account verification became a paid feature. This raises concerns about the potential consequences of monetized verification: by enabling users to purchase visibility and prestige, the platform may open the door to malicious actors seeking to manipulate discourse and distort the public agenda. A deeper comparison between “old” –i.e. pre-November 2022– and “new” –i.e. post-November 2022– verified users will be the target of further research.
Competing interests
The authors declare that they have no competing interests.
Acknowledgements
The authors are thankful to Lorenzo Federico for interesting discussions. SG and FS were partially supported by the project “CODE – Coupling Opinion Dynamics with Epidemics”, funded under PNRR Mission 4 “Education and Research” - Component C2 - Investment 1.1 - Next Generation EU “Fund for National Research Program and Projects of Significant National Interest” PRIN 2022 PNRR, grant code P2022AKRZ9.
Appendix A Maximum entropy null models for bipartite networks
A.1 Bipartite Configuration Model
The framework described in Section 2.1 can be adapted to bipartite networks. In particular, in the following we will use the degree sequences of nodes of both layers as constraints: in this case, the null-model takes the name of Bipartite Configuration Model (or BiCM [40]).
Let us call the two layers of a bipartite network $\top$ and $\bot$ and let $N_{\top}$ and $N_{\bot}$ be their dimensions; nodes belonging to the $\top$-⊤ -layer will be denoted by Latin indices, while the ones belonging to the $\bot$-⊥ -layer will be denoted by Greek ones. Any bipartite network is completely described by its biadjacency matrix $\mathbf{B}$, i.e. a $N_{\top}\times N_{\bot}$ matrix whose entries $b_{i\alpha}=1$ if nodes $i\in\top$ is connected to $\alpha\in\bot$ and $b_{i\alpha}=0$ otherwise. The degree of a node $i\in\top$ ($\alpha\in\bot$), i.e. the number of connections that the nodes have, is simply $k_{i}=\sum_{\alpha}b_{i\alpha}$ ($h_{\alpha}=\sum_{i}b_{i\alpha}$).
Following the procedure sketched in Section 2.1 returns that the probability of observing the generic (bipartite) graph $G_{\text{Bi}}$ is
$\begin{split}P(G_{\text{Bi}})=&\prod_{i,\alpha}\dfrac{e^{-(\theta_{i}+\eta_{\alpha})b_{i\alpha}(G_{\text{Bi}})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}}\end{split}$ (3)
where $\vec{\theta}$ and $\vec{\eta}$ are the Lagrangian multipliers associated with the degree sequences on, respectively, layers $\top$ and $\bot$. Eq. (3) indicates that probabilities per graph factorize in terms of independent probabilities per link $p_{i\alpha}=\frac{e^{-(\theta_{i}+\eta_{\alpha})}}{1+e^{-(\theta_{i}+\eta_{\alpha})}}$. To get the numerical values of the Lagrangian multipliers, we can impose the likelihood maximization [38].
In the present analysis, we used the python module bicm to calculate the numerical values of the probabilities: bicm, described in Ref. [49] as a submodule of NEMtropy, is also a standalone module.
A.1.1 BiCM validation of the monopartite projection of a bipartite network
Using BiCM as a benchmark permits to validate the co-occurrences in the real network [50]. In fact, observed co-occurrences are simply the “projection” of the information contained in the bipartite network into one of the two layers, say $\top$:
$V^{ij}=\sum_{\alpha}b_{i\alpha}b_{j\alpha},\,\forall i\neq j\in\top$
(in Ref. [40] co-occurrences were also called V-motifs since, if the two layers of the bipartite networks are drawn horizontally, they sketch a “V” between the layers). Since in BiCM the probabilities per link are independent, then
$\langle V^{ij}\rangle=\sum_{\alpha}p_{i\alpha}p_{j\alpha},\,\forall i\neq j\in\top.$
In practice, the entire BiCM distribution of each $V^{ij}$ can be calculated: they are Poisson-Binomial distributions, i.e. the extension of Binomial distribution in which each event has a different probability. Calculating exactly the Poisson-Binomial distribution is computationally costly. It was shown, that when probabilities are relatively small –as in our case–, the Poisson-Binomial distribution can be finely approximated with a Poisson distribution [51]. Therefore, it is possible to efficently calculate the p-value associated with each observed co-occurrence $(V^{ij})^{*}$. Since the number of different co-occurrences is $\binom{N_{\top}}{2}$, we need multiple hypothesis testing. In the present manuscript, we implemented the False Discovery Rate (FDR,  [46]) since it controls the rate of False Positives. Again, also the validated projection used in the present manuscript was calculated using the python module bicm.
A.2 Bipartite Directed Configuration Model
Let us now consider the case in which we have a bipartite directed network and we want our model to include the information about the in- and out-degrees of nodes on both layers. It can be shown that the probability per bipartite directed graph $G_{\text{BiD}}$ factorises in terms of the BiCM probability per two undirected bipartite graphs $G_{\text{Bi}}^{\top\rightarrow\bot}$ and $G_{\text{Bi}}^{\top\leftarrow\bot}$, describing, respectively, all edges from layer $\top$ to layer $\bot$, and vice versa [41, 18]:
$P_{\text{BiDCM}}(G_{\text{BiD}})=P_{\text{BiCM}}(G_{\text{Bi}}^{\top\rightarrow\bot})\cdot P_{\text{BiCM}}(G_{\text{Bi}}^{\top\leftarrow\bot}).$
If $b_{i\alpha}^{\top\rightarrow\bot}$ and $b_{i\alpha}^{\top\leftarrow\bot}$ are the generic entries of the biadjacency matrices associated to, respectively, $G_{\text{Bi}}^{\top\rightarrow\bot}$ and $G_{\text{Bi}}^{\top\leftarrow\bot}$, then:
$\begin{split}P_{\text{BiDCM}}(G_{\text{BiD}})=&\Big{[}\prod_{i,\alpha}(q_{i\alpha}^{\top\rightarrow\bot})^{b_{i\alpha}^{\top\rightarrow\bot}}(1-q_{i\alpha}^{\top\rightarrow\bot})^{(1-b_{i\alpha}^{\top\rightarrow\bot}})\Big{]}\\ &\cdot\Big{[}\prod_{i^{\prime},\alpha^{\prime}}(q_{i^{\prime}\alpha^{\prime}}^{\top\leftarrow\bot})^{b_{i\alpha}^{\top\leftarrow\bot}}(1-q_{i\alpha}^{\top\leftarrow\bot})^{(1-b_{i\alpha}^{\top\leftarrow\bot}})\Big{]},\end{split}$
where $q_{i\alpha}^{\top\rightarrow\bot}$ and $q_{i\alpha}^{\top\leftarrow\bot}$ are the (independent) probability to observe a link from $i$ to $\alpha$ or vice versa. Following the same procedure described above for BiCM, $q_{i\alpha}$s can be expressed in terms of the Lagrangian multipliers associated with the (constrained) maximization of the entropy, i.e.
$\begin{split}q_{i\alpha}^{\top\rightarrow\bot}=&\dfrac{e^{-(\theta_{i}^{\text{out}}+\eta_{\alpha}^{\text{in}})}}{1+e^{-(\theta_{i}^{\text{out}}+\eta_{\alpha}^{\text{in}})}};\\ q_{i\alpha}^{\top\leftarrow\bot}=&\dfrac{e^{-(\theta_{i}^{\text{in}}+\eta_{\alpha}^{\text{out}})}}{1+e^{-(\theta_{i}^{\text{in}}+\eta_{\alpha}^{\text{out}})}}.\end{split}$
The numerical values of the $\theta_{i}^{\text{out}}$, $\theta_{i}^{\text{in}}$, $\eta_{\alpha}^{\text{out}}$ and $\eta_{\alpha}^{\text{in}}$ can be obtained through the maximization of the likelihood [38].
A.2.1 BiDCM validation of the monopartite projection of a bipartite network
Following the procedure described in Appendix A.1.1, we can provide a statistical validation of the monopartite multilink directed network obtained by projecting, say, on layer $\top$, a bipartite directed network. The number of edges from node $i$ to node $j$ in the projection is
$\mathcal{V}^{ij}=\sum_{\alpha}b_{i\alpha}^{\top\rightarrow\bot}b_{j\alpha}^{\top\leftarrow\bot}.$
The probability distribution of $\mathcal{V}^{ij}$, following the BiDCM, is again a Poisson-Binomial and, since also in the present case the probabilities are small, the Poisson-Binomial can be finely approximated by a Poisson distribution. Therefore we are interested in calculating $\langle\mathcal{V}^{ij}\rangle$ because it is the only parameter of the Poisson distribution. Since all $b_{i\alpha}^{\top\rightarrow\bot}$ and $b_{i\alpha}^{\top\leftarrow\bot}$ are independent, it is simply
$\langle\mathcal{V}^{ij}\rangle=\sum_{\alpha}q_{i\alpha}^{\top\rightarrow\bot}q_{j\alpha}^{\top\leftarrow\bot}.$ (4)
A.2.2 The application of BiDCM-validated projection to the bipartite network of users and posts
In the analyses above, the BiDCM-validated projection was used to validate the multilink retweet network among users. The original system is a bipartite directed network in which the two layers $\top$ and $\bot$ are, respectively, users and tweets: an arrow goes from a user to a tweet if the user is the author of the tweet and vice versa if the user retweeted the tweet. Since every post has a single author, in $G_{\text{Bi}}^{\top\rightarrow\bot}$ each node $\alpha\in\bot$ has $\kappa_{\alpha}^{\text{in}}=\sum_{i}b_{i\alpha}^{\top\rightarrow\bot}=1$. With such an identification, the probability that the user $i$ is the author of the post $\alpha$ is independent of $\alpha$ and reads:
$q_{i\alpha}^{\top\rightarrow\bot}=\dfrac{e^{-(\theta_{i}^{\text{out}}+\eta^{\text{in}})}}{1+e^{-(\theta_{i}^{\text{out}}+\eta^{\text{in}})}};$
in this case, the Bipartite Configuration Model reduces to a Bipartite Partial Configuration Model [50]. Imposing that the constraints, i.e. the number of tweets authored by each user and the fact that each post has a single author, returns
$q_{i\alpha}^{\top\rightarrow\bot}=\dfrac{\kappa_{i}^{\text{out}}}{N_{\bot}},\quad\forall\alpha\in\bot.$ (5)
Such identification is particularly effective in our case, since plugging Eq. (5) in Eq. (4), gives
$\langle\mathcal{V}^{ij}\rangle=\sum_{\alpha}q_{i\alpha}^{\top\rightarrow\bot}q_{j\alpha}^{\top\leftarrow\bot}=\dfrac{\kappa_{i}^{\text{out}}}{N_{\bot}}\sum_{\alpha}q_{j\alpha}^{\top\leftarrow\bot}=\dfrac{\kappa_{i}^{\text{out}}\kappa_{j}^{\text{in}}}{N_{\bot}},$ (6)
where we used that the degree sequence is constrained in $G_{\text{Bi}}^{\top\leftarrow\bot}$. The result in Eq. (6) is remarkable: the expected value $\langle\mathcal{V}^{ij}\rangle$ can be easily calculated without knowing explicitly the value of the Lagrangian multipliers. Therefore, p-values against BiDCM of the observed $\mathcal{V}^{ij}$ can be easily calculated, and further statistically validated using the FDR procedure described above [46].
A.2.3 Limits of applicability
Besides the simplicity of Eq. (6), there is an important issue, already mentioned for the present application in Ref. [18]. In the null model implemented above, constraints are satisfied on average. Therefore, there are configurations in the ensemble that are not physical, i.e. posts that may have either more than a single author or none. Since both non-physical configurations, i.e. more authors or none, contribute to calculating the p-values in opposite directions, their final effect is hard to predict. Nevertheless, in light of the present results and the ones of previous research, the agreement with the ground truth seems to indicate that, at least in the present application, the final effect of such an approximation is limited. Such an issue is going to be the target of further research.
Appendix B The V-measure for clustering similarity
Let us define, respectively the homogeneity and the completeness of $Q$ with respect to $P$ as
$h(P,Q)$ $=1-\frac{S(P\mid Q)}{S(P)}$
$c(P,Q)$ $=1-\frac{S(Q\mid P)}{S(Q)}.$
Note that $h=1$ if and only if the clusters obtained with $Q$ are sub-clusters of those obtained with $P$, whereas $c=1$ if and only if the clusters obtained with $Q$ are super-clusters of those obtained with $P$. It can be shown that the V-measure $\mathrm{VM}_{\beta}(P,Q)$ is the harmonic mean of $h$ and $c$, weighted by the parameter $\beta$ [31]:
$\mathrm{VM}_{\beta}(P,Q)=\frac{(1+\beta)h(P,Q)c(P,Q)}{\beta h(P,Q)+c(P,Q)}.$
It is worth noting that the VM of two partitions does not decrease linearly with the number of elements on which the two partitions disagree: the $\mathrm{VM}_{1}$ between two partitions obtained with the same algorithm is generally not greater than 0.9; if we consider two partitions that differ by only a small percentage of the network that is randomly relabeled, their $\mathrm{VM}_{1}$ is about 0.8 with $\sim$5of the network relabeled, and decreases to about 0.6 when $\sim$10-13of the network has been relabeled (cfr. Fig. 9 in Appendix B). Based on our tests, we deduce that a reasonably good $\mathrm{VM}_{1}$ between two partitions obtained with different methods/algorithms lies in the range $[0.6-0.9]$.
Fig. 9 shows the average $\mathrm{VM}_{1}$ between some partition of the Elections dataset and a noisy version of the same partition obtained randomly relabeling part of the users. We consider two choices for the initial labeling: 5 equally sized random communities and the communities obtained running the Louvain algorithm. In both cases, the $\mathrm{VM}_{1}$ rapidly decreases to about 0.8 when $\sim$5of the network has been relabeled, and to 0.6 when $\sim$10-13of the network has been relabeled. We also show that two different runs of the Louvain algorithm generate partitions whose $\mathrm{VM}_{1}$ is about 0.9, on average.
Figure 9: Understanding the values of $\mathrm{VM}_{1}$. The plot shows $\mathrm{VM}_{1}$ between independent runs of the Louvain algorithm compared with the $\mathrm{VM}_{1}$ between some partition of the Elections dataset –either 5 equally sized random communities, or the communities obtained running the Louvain algorithm– and a noisy version of the same partition obtained randomly reallocating part of the users, as a function of the percentage of the network that is reallocated.
Appendix C Additional analyses
In the following, we report additional figures and tables that have been left out of the main sections. For the sake of readability and clarity, all discussions can be found in the respective captions.
Figure 10: The average $\mathrm{VM}_{1}$ between partitions obtained with off-the-shelf algorithms and 100 runs of the Louvain algorithm (in green, the self-similarity of Louvain partitions). Louvain was chosen as the benchmark, but all standard methods display comparable performances. Standard algorithms show great variability in the communities assigned to verified users.
Figure 11: The $\mathrm{VM}_{1}$ between partitions of the entire network, obtained with different methods, and the propagated annotations. Let us note that, when comparing the results of the propagated labels of $\mathrm{MonoDC}_{V}$ and $\mathrm{BiDC}$ with the annotations of verified users propagated using the same label propagation algorithm, our methods have an advantage. Nonetheless, the performances of standard methods are worse, even when limited to annotated users.
Table 3: Mean and standard deviation of the execution time, expressed in seconds, of the different steps of $\mathrm{BiDC}_{V}$ on the Elections dataset, the largest dataset considered in this paper. These timings highlight that the cost of $\mathrm{BiDC}_{V}$ is dominated by the label propagation algorithm presented in Ref. [42].
mean std
algorithm step
Label propagation on rt network 943.76 62.14
Louvain on top network 2.64 0.03
validated projection p=0.01 69.74 2.42
Figure 12: An analysis of the role of validation ($\alpha$ denotes the significance level) in determining the partitions obtained for the $\top$ layer of $\mathrm{BiDC}_{V}$. Top: the $\mathrm{VM}_{1}$ between partitions of annotated $V$ users, obtained for different validation levels, and the annotations. Bottom: the $\mathrm{VM}_{1}$ between partitions of $V$ users, obtained for different validation levels, and those obtained with $\alpha=0.01$. Each point is a different run of the method indicated by the color and marker. The results on the non-validated network are very poor, whereas the significance level has a limited impact on the inferred partitions.
Figure 13: The distribution of the H-index in the three datasets. The threshold $H=3$ used to tell apart $\top$ and $\bot$ users is justified by the empirical H-index distribution having an elbow at 3; cumulatively, only about $1-2\%$ of the users pass the threshold.
Figure 14: Cross-comparison of the partitions obtained with $\mathrm{MonoDC}_{\top}$ and $\mathrm{BiDC}_{\top}$, for different choices of $\top$. The $\mathrm{VM}_{1}$ between partitions obtained with different choices of $\top$ and those obtained with $\top=V$, used as a reference, for the Elections dataset. The columns refer to the user set upon which the VM is computed. For instance, “user set = influential” means that in the second column we only consider how influential users have been labeled by different methods, temporarily ignoring the rest of the network.